he controller of Furgee Industries has collected the following monthly expense data for use in analyzing the cost behaviour of maintenance costs.

Month

Total
Maintenance Costs

Total
Machine Hours

January $2,700 230
February 3,350 280
March 4,400 430
April 5,870 620
May 3,700 330
June 6,300 630

(a1)

Determine the variable cost component using the high-low method.
Variable cost per machine hour $
Answer #1

High-low Method:

Variable cost Per Machine hours

= ( Total cost at High machine-hours - Total Cost Low Machine Hours)/ ( High Machine hours - Low Machine Hours)

High Activity = June ( MH =630, Cost = 6,300)

Low Activity = January ( MH 230 , cost = 2,700)

=(6300-2700) / ( 630-230)

=( 3600/400) = $ 9.00 Per MH

Answer : Variable cost Per Machine hour = $ 9.00